OpenStudy (anonymous):

find ther derivative of: f(x)=4/cotx + 6x^2 i have: -4/csc^2(x) + 12x is that right?

OpenStudy (anonymous):

4/cot x= 4 tanx

OpenStudy (anonymous):

and the derivative of 4 tanx = 4 sec^2 x

OpenStudy (anonymous):

do u know the derivative of 1/x?

OpenStudy (anonymous):

i meant sec not csc

OpenStudy (anonymous):

2 root x

OpenStudy (anonymous):

or -x^2

OpenStudy (anonymous):

or somthing

OpenStudy (anonymous):

derivative of 1/x = -1/x^2

OpenStudy (anonymous):

but how did it come?

OpenStudy (anonymous):

okay but the derivative of cot = -csc2x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but u have to find the derivative of 4/cot x

OpenStudy (anonymous):

just wait...derivative of x is 1..right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

den will the derivative of 1/x be also 1?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the derivative of 1/x = -1/x^2

OpenStudy (anonymous):

ok...so derivative of cot x is - cosec^2 x den derivative of 1/cot x is - 1/ cosec^2x?

OpenStudy (anonymous):

r u getting..what do i mean?

OpenStudy (anonymous):

i know all my derivatives and my trig derivatives, like i dont understand what you're trying to say here.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i just want to know why my answer isnt right and what the right answer is.

OpenStudy (anonymous):

right answer is 4 sec^2 x+12 x

OpenStudy (anonymous):

why

OpenStudy (anonymous):

yep confirmed askash123 is right !

OpenStudy (anonymous):

have u ever encountered this d (u/v) / dx= { u' v - u v'}/ v^2 where u and v are functions of x and u' and v' are the derivative of u and v wrt x respectively

OpenStudy (anonymous):

see u have to find the derivative of 4/cotx...right?

OpenStudy (anonymous):

and u know the formula for the derivative of cotx?

OpenStudy (anonymous):

i.e. d(cot x)/ dx= - cosec^2 x

OpenStudy (anonymous):

and u r using this in finding out the derivative of 1/ cotx

OpenStudy (anonymous):

so why i asked u...d(1/x)/ x= ?

OpenStudy (anonymous):

it should be 1 because the derivative of x is 1 so d(1/x)/ dx=1

OpenStudy (anonymous):

the derivative of x is 1 so d(1/x)/ dx=1

OpenStudy (anonymous):

u agree that derivative of 1/x is -1/x^2 but how did it come?